# Prompt Guiding Multi-Scale Adaptive Sparse Representation-driven Network for Low-Dose CT MAR

### v1

@@ V-B Implementation Details and Evaluation Criteria
During the training of PDuMSRNet, we utilized a total of 3000 images, comprising doses of 1/2, 1/4, and 1/8, with each dose containing 1000 images, to train a universal model. Conversely, during the training of DuMSRNet, we trained three distinct models on images with 1/2, 1/4, and 1/8 dose levels, utilizing 1000 CT images for each dose level. Additionally, to further demonstrate the effectiveness of the prompt guidance strategy, similar to PDuMSRNet, we employed 3000 images to train DuMSRNet and obtain a single model, resulting in a blind DuMSRNet (BDuMSRNet).
@@
V-C Performance Evaluation for LDMAR Task
To demonstrate the superiority of DuMSRNet, we conduct both quantitative and qualitative comparisons with previous LDCT methods, such as LWFSN (Zavala-Mondragon et al., 2022), CNN10 (Ghani and Karl, 2018), REDCNN (Chen et al., 2017), CTformer (Wang et al., 2023a), EDCNN (Liang et al., 2020), OctaveNet (Won et al., 2020) and InDuDoNet (Wang et al., 2021), under different dose levels. All methods are trained and tested on the same dataset. Table I summarizes the quantitative results of different methods. It is evident that DuMSRNet achieves the best performance across all metrics under both dose levels. Fig. 3 illustrates the visual comparison between DuMSRNet and other SOTA methods under the 1/2 dose level. As observed, the FBP reconstructed images are significantly affected by noise and metal artifacts. While LWFSN, CNN10, REDCNN, CTformer, EDCNN, and OctaveNet mitigate certain metal artifacts, they concurrently introduce new noise and metal artifacts. The MAR method InDuDoNet successfully remove most metal artifacts, but they miss tissue details. By contrast, DuMSRNet effectively removes most metal artifacts while preserving the tissue structures in CT images, which demonstrates the importance of multi-scale and within-scale information for LDMAR task. Fig. 4 presents the reconstruction results of various methods under the 1/4 dose level. As shown in Fig. 4, even with a further reduction in dose level to 1/4, DuMSRNet can still remove more noise and artifacts, achieving optimal reconstruction performance.
To validate the efficacy of our prompt guiding strategy, we devised two models: the blind DuMSRNet, trained on a dataset comprising three doses without the prompt guiding strategy, and the PDuMSRNet, trained on a dataset with three doses using the prompt guiding strategy. As shown in Table I, we observed a notable improvement of 0.34 dB in PSNR value for the PDuMSRNet compared to that of the Blind DuMSRNet. This substantial enhancement underscores the efficacy of the prompt guiding strategy. Leveraging prompt information enables the proposed DuMSRNet to better comprehend the intricate relationship between dose and noise, as well as the nuanced mapping between metal masks and artifacts, thereby achieving superior reconstructions.
@@ V-C Performance Evaluation for LDMAR Task
Figure 5: Visual evaluation and reconstruction error magnitude of BDuMSRNet and PDuMSRNet on LDCT images under the 1/4 dose level. Zooming in on the green screen enhances the viewing experience of the images. The PSNR and SSIM values are computed on the entire image reconstructed by the respective approaches. The metallic implants are highlighted with red masks.
Fig. 5 illustrates the reconstruction outcomes obtained by BDuMSRNet and PDuMSRNet under the 1/4 dose level. As depicted, DuMSRNet exhibits superior noise and artifacts removal capabilities compared to BDuMSRNet, resulting in enhanced reconstruction performance. This improvement highlights the effectiveness of the prompt embedding strategy.
TABLE II: Storage requirement for different models. In comparison regarding the storage costs (KB) of models for three doses, the multiple-in-one model demonstrates better storage efficiency
Single-dose models Muti-doses models
Method DuMSRNet BDuMSRNet PDuMSRNet
Storage 158,856 52,952 52,996
V-D Performance Evaluation on Storage Requirement
To demonstrate the storage efficiency superiority of DuMSRNet, we conduct a qualitative comparison of model storage requirements with the traditional single-dose models (DuMSRNet) and the multi-dose models (BDuMSRNet and PDuMSRNet). As illustrated in Table II, DuMSRNet incurs the highest storage costs among the models. The storage requirement of PDuMSRNet is approximately equivalent to BDuMSRNet, which is one-third that of DuMSRNet. Although the storage cost of the PDuMSRNet is slightly higher than that of the BDuMSRNet, the reconstruction quality achieved by the PDuMSRNet is better than that achieved by the BDuMSRNet.
@@ V-D Performance Evaluation on Storage Requirement
TABLE III: Ablation studies of the dual domain learning under the 1/2 dose level. Best values are highlighted in bold.
@@
V-E Ablation Studies
(1) The impact of dual domain learning: During each unfolding stage, DuMSRNet includes a network for the sinogram domain and a network for the image domain. To showcase the effect of dual domain learning, we train sinogram-domain-only and image-domain-only networks under identical conditions. We then compare their performance with DuMSRNet using PSNR, SSIM, and RMSE metrics. As summarized in Table III, it is evident that our dual domain approach (i.e., DuMSRNet) outperforms the sinogram-domain-only and image-domain-only methods, demonstrating the effectiveness of leveraging joint dual domain information.
TABLE IV: Ablation studies on different modules of DuMSRNet under the 1/2 dose level. We report the PSNR (dB) / SSIM / RMSE (HU) values of the testing dataset for each case. Best values are highlighted in bold.
@@ V-E Ablation Studies
we employ a learnable parameter as the threshold, replacing the threshold generated by SATG, forming DuMSRNet w/o SATG. Note that SATG here does not contain the prompt guiding module. When combined with DuMSRNet and DuMSRNet w/o SATG, the proposed SATG produces a significant performance improvement on all metrics, as summarized in Table IV. To assess the effectiveness of the combination of the deep feature extraction module with three branches and the feature selection module in SATG, we selectively remove the corresponding parts during training, resulting in DuMSRNet w/o LIEB, DuMSRNet w/o RIEB, DuMSRNet w/o GIEB, and DuMSRNet w/o FEM. The corresponding quantitative results are summarized in Table IV. The comparison among DuMSRNet w/o LIEB, DuMSRNet w/o RIEB, DuMSRNet w/o GIEB, and DuMSRNet validates that the combination of local, regional, and global information significantly enhances the reconstruction performance. Additionally, compared to DuMSRNet w/o FEM, DuMSRNet achieves better performance, thanks to the capability of the feature selection module. Specifically, the feature selection module can reweight local, regional, and global information, preventing performance degradation caused by information redundancy.
TABLE V: Ablation studies on different prompts of PDuMSRNet under the 1/2, 1/4 and 1/8 dose levels. We report the PSNR (dB) / SSIM / RMSE (HU) values of the testing dataset for each case. Best values are highlighted in bold.
@@
(3) The impact of prompt guiding strategy: The second group pertains to PDuMSRNet models across multiple dose levels (1/2, 1/4, and 1/8) to validate the effectiveness of the prompt guiding strategy. To facilitate the adaptation of a single model to various CT dose settings, we introduce a prompt guiding strategy that incorporates the prompt information from input LDCT images with metallic implants, mental masks, and dose maps to guide the training process. To evaluate the effectiveness of different prompts, we systematically remove input LDCT images with metallic implants, mental masks, and dose maps, and replace the 1/dose map with the dose map in separate instances. As illustrated in Table V, the performance in these ablation experiments falls short of that of the original PDuMSRNet, thereby confirming the validity of each individual prompt.
TABLE VI: Ablation studies on MSFuM of DuMSRNet under the 1/2 dose level. We report the PSNR (dB) / SSIM / RMSE (HU) values of the testing dataset for each case. Best values are highlighted in bold.
@@
(4) The impact of fusing high-resolution and low-resolution information: To fully utilize cross-scale complementarity, we design a multi-scale coefficient fusion module using the cross attention mechanism, achieving efficient fusion of multi-scale information. To validate the effectiveness of fusing high-resolution and low-resolution information, we modify the original network architecture to create DuMSRNet w/o MSFuM, where the architecture no longer performs the fusion of high-resolution and low-resolution information. As shown in Table VI, the fusion mechanism of high-resolution and low-resolution leads to an improvement in PSNR values by 0.89 dB, SSIM values by 0.0086, and a reduction in RMSE values by 2.28.
VI Conclusion
@@
In this paper, we proposed a model interpretable prompt guiding multi-scale adaptive sparse representation-driven network named PMSRNet for LDMAR, and further constructed PDuMSRNet, an enhancement of DuMSRNet in dual domain. Specifically, the PMSRNet consists of three components: multi-scale sparsifying frames, PSATG and MSFuM. The proposed multi-scale sparsifying frames could simultaneously employ within-scale characteristics and cross-scale complementarity. The elaborated PSATG could extract features at the local, regional, and global range to capture multiple contextual information. Additionally, the introduced MSFuM could fuse detailed information from high-resolution features and contextual information from low-resolution features. Extensive experiments demonstrated that the proposed DuMSRNet consistently outperformed the state-of-the-art methods and the proposed PDuMSRNet achieved less model storage cost and more flexibility.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants 62371414 and 61901406, by the Natural Science Foundation of Hebei Province under Grant F2023203043 and F2020203025, by the Young Talent Program of Universities and Colleges in Hebei Province under Grant BJ2021044, and by the Hebei Key Laboratory Project under Grant 202250701010046.

### v3

@@ 5.3 Implementation Details and Evaluation Criteria
We design three training strategies to evaluate performance of the proposed algorithm for LDMAR task, DuMSRNet, PDuMSRNet and blind DuMSRNet, i.e., BDuMSRNet. Specifically, DuMSRNet is trained at single-dose level to compare with other single-dose SOTA methods. In contrast, PDuMSRNet and BDuMSRNet are trained across all three dose levels. BDuMSRNet is the vision removing prompt guidance strategy compared with PDuMSRNet to valid the effect of prompt guidance strategy. During the training of PDuMSRNet, we utilize a total of 3000 images, comprising doses of 1/2, 1/4, and 1/8, with each dose containing 1000 images, to train a universal model. Conversely, during the training of DuMSRNet, we train three distinct models on images with 1/2, 1/4, and 1/8 dose levels, utilizing 1000 CT images for each dose level. Additionally, to further demonstrate the effectiveness of the prompt guidance strategy, similar to PDuMSRNet, we employ 3000 images to train DuMSRNet and obtain a single model, resulting in BDuMSRNet.
@@
Fig. 3: Visual evaluation of different approaches on LDCT images with metallic implants under the 1/2 dose level. Zooming in on the green screen enhances the viewing experience of the images. The PSNR and SSIM values are computed on the entire image reconstructed by the respective approaches. The metallic implants are highlighted with red masks. The blue arrow indicates the reconstructed structural information. The display window is [-1000, 1000] HU.
Fig. 4: Visual evaluation of different approaches on LDCT images with metallic implants under the 1/4 dose level. Zooming in on the green screen enhances the viewing experience of the images. The PSNR and SSIM values are computed on the entire image reconstructed by the respective approaches. The metallic implants are highlighted with red masks. The blue arrow indicates the reconstructed structural information. The display window is [-1000, 1000] HU.
5.4 Performance Evaluation for LDMAR Task
To demonstrate the superiority of DuMSRNet, we conduct both quantitative and qualitative comparisons with previous LDCT methods, such as FBP, LWFSN (9721076, 9721076), REDCNN (7947200, 7947200), CNN10 (5401721, 5401721), EDCNN (9320928, 9320928), OctaveNet (10.1007/978-3-030-59354-4_7, 10.1007/978-3-030-59354-4_7), CTformer (Wang_2023, Wang_2023), LIT-Former (Chen2023LITFormerLI, Chen2023LITFormerLI), InDuDoNet (10.1007/978-3-030-87231-1_11, 10.1007/978-3-030-87231-1_11), CLIPDenoising (Cheng_2024_CVPR, Cheng_2024_CVPR), and multi-task learning approaches, such as AirNet (airnet, airnet), PromptIR (promptir, promptir). under different dose levels. All methods are trained and tested on the same dataset. Table 1 reports the average PSNR, SSIM and RMSE values for the DeepLesion dataset at $1/2$, $1/4$ and $1/8$ dose levels. Table 2 summarizes the average PSNR and SSIM values on varying metal sizes under $1/8$ dose level. As can be seen from Table 1 and Table 2, all DL-based approaches outperform traditional methods in terms of PSNR and SSIM values, owing to the superior feature extraction and representation capabilities inherent to deep neural networks. Notably, the dual-domain frameworks, InDuDoNet and our DuMSRNet, achieve higher PSNR values compared to single-domain methods, underscoring the benefit of leveraging knowledge from both the sinogram and image domains. Our proposed PMDeepST demonstrates consistent and significant superiority across all evaluated dose levels and metal sizes. Fig. 3 and Fig. 4 illustrate the visual comparison between DuMSRNet and other SOTA methods under the 1/2 and 1/4 dose level and the visual comparisons critically illuminate the reasons behind our quantitative advantages. As observed, the FBP reconstructed images are severely degraded by noise and metal artifacts. While LWFSN, CNN10, REDCNN, CTformer, EDCNN, and OctaveNet achieve some artifact suppression, they concurrently introduce secondary noise and metal artifacts. InDuDoNet, though effective in removing most metal artifacts, tends to over-smooth the results, leading to a loss of fine anatomical details, which is associated with its ignoring on multi-scale feature within the CT image. By contrast, DuMSRNet more effectively suppresses a wider range of artifacts while simultaneously preserving critical tissue structures. This performance is directly linked to the core architectural strength of our method, the multi-scale adaptive sparse representation framework and the prompt guiding module, which is explicitly designed to model within-scale characteristics, cross-scale complementarity and the prior of metal implants. This design enables the network to distinguish between artifact patterns and true anatomical textures more reliably than methods relying on a single-scale or less adaptive representations. However, a critical analysis also reveals a limitation when processing larger metal implants. In these challenging cases, the network occasionally fails to distinguish massive streaking artifacts from true anatomical tissues, indicating that the current multi-scale sparsifying frames struggle to effectively model and remove large-scale metal artifacts. This observed shortcoming points toward a direction for future work aimed at enhancing the representational capacity for such extreme scenarios.
To validate the efficacy of our prompt guiding strategy, we devised two models: the blind DuMSRNet, trained on a dataset comprising three doses without the prompt guiding strategy, and the PDuMSRNet, trained on a dataset with three doses using the prompt guiding strategy. As shown in Table 1, we observed a notable improvement of 0.39 dB in PSNR value for the PDuMSRNet compared to that of the Blind DuMSRNet. This substantial enhancement underscores the efficacy of the prompt guiding strategy. Leveraging prompt information enables the proposed DuMSRNet to better comprehend the intricate relationship between dose and noise, as well as the nuanced mapping between metal masks and artifacts, thereby achieving superior reconstructions. Fig. 5 illustrates the reconstruction outcomes obtained by BDuMSRNet and PDuMSRNet under the 1/4 dose level. As depicted, DuMSRNet exhibits superior noise and artifacts removal capabilities compared to BDuMSRNet, resulting in enhanced reconstruction performance. This improvement highlights the effectiveness of the prompt-embedding strategy. Moreover, since PDuMSRNet can achieve image reconstruction under three dose levels with a single model, it requires less model storage and offers greater flexibility than DuMSRNet.
Fig. 5: Visual evaluation and reconstruction error magnitude of BDuMSRNet and PDuMSRNet on LDCT images under the 1/4 dose level. Zooming in on the green screen enhances the viewing experience of the images. The PSNR and SSIM values are computed on the entire image reconstructed by the respective approaches. The metallic implants are highlighted with red masks.
@@ 5.4 Performance Evaluation for LDMAR Task
Table 3: Ablation studies of the dual domain learning under the 1/2 dose level. Best values are highlighted in bold.
@@
5.5 Ablation Studies
(1) The impact of dual domain learning: During each unfolding stage, DuMSRNet includes a network for the sinogram domain and a network for the image domain. To showcase the effect of dual domain learning, we train sinogram-domain-only and image-domain-only networks under identical conditions. We then compare their performance with DuMSRNet using PSNR, SSIM, and RMSE metrics. As summarized in Table 3, it is evident that our dual domain approach (i.e., DuMSRNet) outperforms the sinogram-domain-only and image-domain-only methods, demonstrating the effectiveness of leveraging joint dual domain information.
Table 4: Ablation studies on different modules of DuMSRNet under the 1/2 dose level. We report the PSNR (dB) / SSIM / RMSE (HU) values of the testing dataset for each case. Best values are highlighted in bold.
@@ 5.5 Ablation Studies
We employ a learnable parameter as the threshold, replacing the threshold generated by SATG, forming DuMSRNet w/o SATG. Note that SATG here does not contain the prompt guiding module. When combined with DuMSRNet and DuMSRNet w/o SATG, the proposed SATG produces a significant performance improvement on all metrics, as summarized in Table 4. To assess the effectiveness of the combination of the deep feature extraction module with three branches and the feature selection module in SATG, we selectively remove the corresponding parts during training, resulting in DuMSRNet w/o LIEB, DuMSRNet w/o RIEB, DuMSRNet w/o GIEB, and DuMSRNet w/o FEM. The corresponding quantitative results are summarized in Table 4. The comparison among DuMSRNet w/o LIEB, DuMSRNet w/o RIEB, DuMSRNet w/o GIEB, and DuMSRNet validates that the combination of local, regional, and global information significantly enhances the reconstruction performance. Additionally, compared to DuMSRNet w/o FEM, DuMSRNet achieves better performance, thanks to the capability of the feature selection module. Specifically, the feature selection module can reweight local, regional, and global information, preventing performance degradation caused by information redundancy.
Table 5: Ablation studies on different prompts of PDuMSRNet under the 1/2, 1/4 and 1/8 dose levels. We report the PSNR (dB) / SSIM / RMSE (HU) values of the testing dataset for each case. Best values are highlighted in bold.
@@
(3) The impact of prompt guiding strategy: The second group pertains to PDuMSRNet models across multiple dose levels (1/2, 1/4, and 1/8) to validate the effectiveness of the prompt guiding strategy. To facilitate the adaptation of a single model to various CT dose settings, we introduce a prompt guiding strategy that incorporates the prompt information from input LDCT images with metallic implants, mental masks, and dose maps to guide the training process. To evaluate the effectiveness of different prompts, we systematically remove input LDCT images with metallic implants, mental masks, and dose maps, and replace the 1/dose map with the dose map in separate instances. As illustrated in Table 5, the performance in these ablation experiments falls short of that of the original PDuMSRNet, thereby confirming the validity of each individual prompt.
Table 6: Ablation studies on MSFuM of DuMSRNet under the 1/2 dose level. We report the PSNR (dB) / SSIM / RMSE (HU) values of the testing dataset for each case. Best values are highlighted in bold.
@@
(4) The impact of fusing high-resolution and low-resolution information: To fully utilize cross-scale complementarity, we design a multi-scale coefficient fusion module using the cross attention mechanism, achieving efficient fusion of multi-scale information. To validate the effectiveness of fusing high-resolution and low-resolution information, we modify the original network architecture to create DuMSRNet w/o MSFuM, where the architecture no longer performs the fusion of high-resolution and low-resolution information. As shown in Table 6, the fusion mechanism of high-resolution and low-resolution leads to an improvement in PSNR values by 0.89 dB, SSIM values by 0.0086, and a reduction in RMSE values by 2.28.
Fig. 6: Visual evaluation of different approaches on clinical dataset with metallic implants. The magnified details are shown in green box. The display window is [-175, 275] HU. Blue arrows indicate key areas, such as tissues near the metallic artifacts, which highlight differences in reconstruction quality among all methods.
Fig. 7: SAM segmentation results for LDCT images and results of our proposed method. Different colors represent distinct segmented regions. Blue arrows indicate key areas, such as tissues near the metallic implants and artifacts, which highlight differences in segmentation quality.
6 Discussion
6.1 Analysis on Storage Requirement
To demonstrate the superior storage efficiency of DuMSRNet, we conduct a qualitative comparison of model storage requirements with the traditional single-dose models (DuMSRNet) and the multi-dose models (BDuMSRNet and PDuMSRNet). As illustrated in Table 7, DuMSRNet incurs the highest storage costs among the models. The storage requirement of PDuMSRNet is approximately equivalent to BDuMSRNet, which is one third of DuMSRNet trained model for each dose level. Although the storage cost of the PDuMSRNet is slightly higher than that of the BDuMSRNet, the reconstruction quality achieved by the PDuMSRNet is better than that achieved by the BDuMSRNet. This performance gain is attributable to the inherent capacity of the multi-task, multi-setting architecture of PDuMSRNet, which leverages shared representations while adapting to specific task conditions through its prompt-guided modules.
Table 7: Storage requirement for different models. In comparison regarding the storage costs (KB) of models for three doses, the multiple-in-one model demonstrates better storage efficiency
Single-dose models Muti-doses models
Algorithms DuMSRNet BDuMSRNet PDuMSRNet
Storage 158,856 52,952 52,996
6.2 Analysis on Clinical Validation
To further validate the clinical applicability of our method, we conducted validation on the clinical dataset, obtained from the public CTPelvic1K library  [^1] CTPelvic1K: https://github.com/MIRACLE-Center/CTPelvic1K. CTPelvic1K dataset is a publicly accessible resource specifically designed for pelvic fracture segmentation. The dataset includes 14 testing volumes labeled with multi-bone, i.e., sacrum, left hip, right hip, and lumbar spine. We evaluate the clinical performance of our model on the CTPelvic1K dataset, employing weights pre-trained on the simulated DeepLesion dataset to assess its generalization capability. Fig. 6 presents visualization evaluation of different approaches on clinical dataset with metallic implants.
As illustrated in Fig. 6, due to low dose and the presence of metallic implants, significant noise and artifacts are observed around the metal regions and algorithms exhibit different reconstruction performance from simulated dataset. Specifically, methods such as OctaveNet, LIT-former and CLIPDenoising exhibit severe distortion and highlight clipping in their reconstruction results, significantly harming diagnostic utility. While LWFSN, REDCNN, CNN10 and InDuDoNet demonstrate more structural coherence, they fail to adequately suppress metal artifacts. CTformer and PromptIR introduce secondary textural artifact while they achieve the most effective metal artifacts reduction. Although AirNet demonstrates satisfactory performance in metal artifact reduction and denoising, its output exhibits significant blurring in critical edge regions, as indicated by the blue arrows in Fig. 6. This degradation of edge detail represents a notable limitation for applications requiring precise structural delineation. In contrast, the proposed PDuMSRNet effectively addresses this weakness. It achieves optimal performance by successfully preserving crucial edge information while simultaneously maintaining high efficacy in both artifact suppression and noise reduction.
6.3 Analysis on Segmentation Tasks
To demonstrate the clinical significance of our proposed method PDuMSRNet, we evaluated its impact on downstream segmentation tasks. Specifically, the Segment Anything Model (SAM) (sam, sam), a SOTA segmentation framework, was deployed using its official implementation. Both original LDCT images with metallic implants and our reconstructed results were processed by the identical SAM instance. As illustrated in Fig. 7, our reconstruction results enabled the segmentation of significantly more discernible anatomical structures compared to the LDCT inputs with metallic implants. This quantitative improvement in segmentable regions directly enhances the clinical utility of the images, validating efficacy of our method in facilitating critical downstream applications such as diagnostic analysis and treatment planning.
7 Conclusion
@@
In this paper, we proposed a model interpretable prompt guiding multi-scale adaptive sparse representation-driven network named PMSRNet for LDMAR, and further constructed PDuMSRNet, an enhancement of DuMSRNet in dual domain. Specifically, the PMSRNet consists of three components: multi-scale sparsifying frames, PSATG and MSFuM. The proposed multi-scale sparsifying frames could simultaneously employ within-scale characteristics and cross-scale complementarity. The elaborated PSATG could extract features at the local, regional, and global range to capture multiple contextual information. Additionally, the introduced MSFuM could fuse detailed information from high-resolution features and contextual information from low-resolution features. Extensive experiments demonstrated that the proposed DuMSRNet consistently outperformed the state-of-the-art methods and the proposed PDuMSRNet achieved less model storage cost and more flexibility. However, current methods exhibit blurring in clinical dataset. This limitation is fundamentally attributable to insufficient model generalization capability and the inherent domain gap between simulated training data and clinical test data. Future work will involve training the model using real-world clinical datasets to enhance reconstruction performance.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants 62371414, by the Natural Science Foundation of Hebei Province under Grant F2025203070, and by the Hebei Key Laboratory Project under Grant 202250701010046.